OpenStudy (anonymous):

Simplify. 6/5+6/10 divided by 8/11-8/2

OpenStudy (anonymous):

take lcm of 5,10 and 11, 2 \[((6(2)+6)/10)/ ((8(2) - 8(11))/22) \]

TheSmartOne (thesmartone):

\[\frac{ \frac{ 6 }{ 5}+\frac{ 6 }{ 10 } }{\frac{ 8 }{ 11 }-\frac{ 8 }{ 2 } }=\frac{ \frac{ 12 }{ 10}+\frac{ 6 }{ 10 } }{\frac{ 16 }{ 22 }-\frac{ 88 }{ 22 } }\]

TheSmartOne (thesmartone):

\[\frac{ \frac{ 18 }{ 10 } }{\frac{ -72}{ 22 } }=\frac{ 18 }{ 10 }\times \frac{ 22 }{ -72 }\]

TheSmartOne (thesmartone):

\[\frac{ 18 }{ 10 }\times \frac{ 22 }{ -72 }=\frac{ 1 }{ 10 }\times \frac{ 22 }{ -4 }=\frac{ 1 }{ 5 }\times \frac{ 11 }{ -4 }=\frac{ 11 }{ -20 }\]

OpenStudy (xxdragoonxx):

6/5+6/10/8/11-8/2 6/5-8/2+6/10/8/11 = 8/2=4 = 6/5-4+6/10/8/11 6/10/8/11 6/10=3/5 = 3/5/8/11 = 3/5*11/8 = 5*8=40 = 3*11/40 = 3*11=33 = 33/40 = 6/5-4+33/40 33-4+6/5=1*33+1*2^3*6-2^3*4*5/40 = 33/40-4+6/5 4=4/1 = -4/1+6/5+33/40 40=2^3*5 1*2^3*5=40 = -2^3*4*5/40+1*33/40+1*2^3*6/40 = 1*33+1*2^3*6-2^3*4*5/40 = -2^3*14+33/40 = -79/40